OpenStudy (anonymous):

Given the expression 5a^2b – 13ab + 7a^3 – 4b, do the following as instructed below: a)Write the polynomial in descending order. b)Classify the polynomial by the number of terms. c)State the degree of the polynomial.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

a) 7a^3 +5a^2b -13ab -4b

OpenStudy (anonymous):

b) Although the ^2b is unusual, the highest power seems to be ^3, therefore it would be a third degree equation.

OpenStudy (anonymous):

sorry, the answer I gave to b should have been the answer to c)

OpenStudy (anonymous):

oh ok but can you explian it to me to

OpenStudy (anonymous):

b) there are 4 terms in the equation, and since we only label equations up to third degree, then it must be called a polynomial

OpenStudy (anonymous):

what would you like me to explain @CUTEYPIE137

OpenStudy (anonymous):

In fact, it says its a polynomial in the question

OpenStudy (anonymous):

well um the descending order how did you get that @BPDlkeme234

OpenStudy (anonymous):

A linear equation is one in which the highest power is 1 (e.g x+ y = 12) A quadratic equation is one in which the highest power is 2 (e.g x^2 + y = 12) a cubic equation is one in which the highest power is 3 (e.g. (x^3 + y^2+3 =12) a polynomial is any equation in which the power is above this

OpenStudy (anonymous):

You asked about the descending order. I simply looked at the powers. It seems to me that the highest power was ^3 so that term came first. next ^2b (as I said this is unusual) but I put it second since we dont know what b is. Next 13ab because it is simply to the power of 1 (i.e. any number to the power of 1 is itself e.g. 13 ^1 = 13), but it had two variables a and b. and finally the last term which had only one variable (-4b) so it went last.

OpenStudy (anonymous):

oh thank you i understand um can you explain question b for me also please @BPDlkeme234

OpenStudy (anonymous):

OH NEVERMIND

OpenStudy (anonymous):

I know it would be a polynomial right @BPDlkeme234

OpenStudy (anonymous):

@CUTEYPIE137 it is asking you to COUNT the number of terms in the expression first term: 7a^3 2nd term: 5a^2b 3rd term: 13ab 4th term: -4b therefore there are 4 terms in this expression.

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

You understand now?@CUTEYPIE137

OpenStudy (anonymous):

so it would be a cubic polynomial @BPDlkeme234

OpenStudy (anonymous):

only if the highest power was ^3 in the expression.

OpenStudy (triciaal):

is everything clear

OpenStudy (anonymous):

yes so in this polynomial the highest power is 3 so it is a cubic polynomial

OpenStudy (anonymous):

Well we dont use the words polynomial if it is a cubic equation, we simply say it is a cubic equation. polynomials are equations in which the power is greater than ^3

OpenStudy (anonymous):

you are right it is a cubic equation, if we dont know what ^2b is.

OpenStudy (anonymous):

oh ok thank so very much @BPDlkeme234

OpenStudy (triciaal):

not an equation

OpenStudy (anonymous):

correct @triciaal I had mentioned previously it was an expression not an equation